a) Complete the table to determine the balance A for P dollars invested at rate r for t years, compounded n times per year. (Round your answers to two decimal places.)

P = $3500, r = 3.5%, t = 20 years

n A
1 $
2 $
4 $
12 $
365 $
Continuous
compounding
$

B)Complete the table to determine the amount of money P that should be invested at rate r to produce a final balance of $120,000 in t years. (Round your answers to two decimal places.)

r = 3%, compounded continuously

t P
1 $
10 $
20 $
30 $
40 $
50 $

c)

Find the effective rate of interest corresponding to a nominal rate of 8.5% per year compounded annually, semiannually, quarterly and monthly. (Round your answers to two decimal places.)

annually__ %

semiannually__ %

quarterly___%

monthly__ %
Answer #1

las P=$3500 8= 3-51. t = 20 years Here nal → A = ? na 2 > A= ? n=4 → A = ? n=12 -A ? n=365 → A-? nt Amount [X20 nel -> A- P(ib) A = $ 120,000 r=3% Compaunded Continuously ta! pa ? t-10 → P= ? t = 20 = P = ? t 30 < P-? E = eo > P = ? t= 50 p= ? (Compoc) The general formula for interest Veft (1+), given r = 8.5 = and for quarterly we compute 14 Wege = (1+0.985) Ñ 0.0877 whic